The current price of Estelle Corporation stock is $ 23.00 In each of the next two​ years, this stock price will either go up by 18 % or go down by 18 %. The stock pays no dividends. The​ one-year risk-free interest rate is 8.0 % and will remain constant. Using the Binomial​ Model, calculate the price of a​ one-year call option on Estelle stock with a strike price of $ 23.00
Answer #1
High price=Current price*up move=23*1.18=27.14
Low price=Current price*down move=23*0.82=18.86
Risk neutral probability for up move
q = (e^(risk free rate*time)-D)/(U-D)
=(e^(0.08*1)-0.82)/(1.18-0.82)=0.73135
Call option payoff at high price (payoff H)
=Max(High price-strike price,0)
=Max(27.14-23,0)
=Max(4.14,0)
=4.14
Call option payoff at low price (Payoff L)
=Max(Low price-strike price,0)
=Max(18.86-23,0)
=Max(-4.14,0)
=0
Price of call option = e^(-r*t)*(q*Payoff H+(1-q)*Payoff L)
=e^(-0.08*1)*(0.731353*4.14+(1-0.731353)*0)
=2.8